OpenStudy (anonymous):

Perform the indicated operation and simplify. (7/5x)-(4/3x)

hero (hero):

1. Factor out x since it's common to both: \[\left(\frac{7}{5} - \frac{4}{3}\right)x\] 2. Subtract the fractions 3. Then simplify

OpenStudy (anonymous):

Out teacher wants us to find the LCD first

hero (hero):

Factoring out the x is a good pre-step to the first step of Finding the LCD.

hero (hero):

Consider factoring out the x "step zero"

OpenStudy (anonymous):

oh okay

hero (hero):

Do you know what the LCD is?

OpenStudy (anonymous):

is it 3x and 5x?

OpenStudy (anonymous):

so 15x?

hero (hero):

Wait, is the original problem: \[\frac{7}{5x} - \frac{4}{3x}\]

OpenStudy (anonymous):

yes

hero (hero):

I see. Well, you can still factor out 1/x: \[\left(\frac{7}{5} - \frac{4}{3}\right)\frac{1}{x}\]

hero (hero):

Now subtract \[\frac{7}{5} - \frac{4}{3}\]

OpenStudy (anonymous):

1/10?

OpenStudy (anonymous):

wait no hold on

hero (hero):

What ever happened to finding the LCD?

OpenStudy (anonymous):

i messed up, just give me a minute

OpenStudy (anonymous):

is it 1/15x, the lcd is 15x?

hero (hero):

I thought we were just subtracting 7/5 - 4/3

OpenStudy (anonymous):

its 1/15

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks :)